solve t1/2+3t1/4+2=0
Answer by checkley77(12844)

t1/2+3t1/4+2=0
t/2+3t/4=-2
(2t+3t)/4=-2
5t/4=-2
5t=-2*4
5t=-8
t=-8/5 ans.
Proof:
-8/5*1/2+3*-8/5*1/4+2=0
-8/10-24/20+2=0
-16/20-24/20+2=0
-40/20+2=0
-2+2=0
0=0